Solve (x+6)1/2= -x

1
Expert's answer
2021-03-19T07:46:56-0400

"\qquad\qquad

\begin{aligned}

\sqrt{x+6}&= -x\qquad\\
(\sqrt{x+6})^2&=(-x)^2\\
(x+6)&=x^2\\
x^2-x-6&=0\\
(x-3)(x+2)&=0\\
\\
x-3=0\qquad&{or}\qquad{x+2=0}\\
x=3\qquad&{or}\qquad{x=- 2}\\\\
When \qquad{x=3};\qquad{\small LHS=\sqrt{9}=+3}\qquad &\small RHS=-3\\
 \small LHS\neq RHS\\
x=-2;\qquad{\small LHS=\sqrt{4}=+2}\qquad& \small RHS=2\\
\small LHS=RHS\\\\
\therefore\qquad \bold{x=-2}
\end{aligned}"